OpenStudy (anonymous):

it says solve the salloing equations write the variables above their valures to find the name of a famous volcano? 4e-8=32?

OpenStudy (anonymous):

e = 10

OpenStudy (anonymous):

show steps ??

OpenStudy (anonymous):

4e -8 = 32 4e = 32+8 4e = 40 e = 40/4 e = 10

OpenStudy (anonymous):

it says solve the salloing equations write the variables above their valures to find the name of a famous volcano a+18= -32??